(v) A lending library has a fixed charge for the first three days and an additional charge for each day thereafter. Saritha paid Rs.27 for a book kept for seven days, while Susy paid Rs.21 for the book she kept for five days. Find the fixed charge and the charge for each extra day.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Let the fixed charge for the first three days be Rs.A and the charge for each day extra be Rs.B. According to the information given, A + 4B = 27 …………………………………….…………………………. (i) A + 2B = 21 ……………………………………………………………….. (ii) When equation (ii) is subtracted from equation (i) we get, 2B = 6 B = 3 …………………………………………………………………………(iii) Substituting B = 3 in equation (i) we get, A + 12 = 27 A = 15 Hence, the fixed charge is Rs.15 And the Charge per day is Rs.3